# Single-Domain Generalized Object Detection by Balancing Domain Diversity and Invariance

### v1

@@ 3.2.1 Diversity Learning Module (DLM)
Traditional domain alignment methods primarily focus on learning domain-invariant features, overlooking the importance of domain-specific information. This limitation restricts the model’s ability to achieve feature diversity, impacting feature extraction. To address this, we introduce the Diversity Learning Module (DLM), which restricts the domain-specific feature to achieve feature diversity. The DLM operates as follows: first, to obtain domain-specific features that capture domain information, we draw inspiration from OCR [17], assuming the relationship between the original and augmented features as $z_{a}=\lambda_{1}z_{s}+(1-\lambda_{1})z_{d}$. Here, $\lambda_{1}$ is a gradually increasing hyperparameter, reflecting the model’s improving capacity to learn domain-invariant features over the process of training. As this capacity grows, the domain-specific information in $z_{s}$ and $z_{a}$ correspondingly diminishes. Ultimately, this approach enables us to obtain the domain-specific feature $z_{d}$ during training.
@@
where $C$ represents the number of categories, $f(z_{d})\in R^{K\times C}$ is prediction function for $z_{d}$. Inspired by [17], to estimate the category semantic information in $z_{d}$, we implement to maximization of the entropy of the prediction results from $z_{d}$.
@@
where $\mathcal{H}$ represents entropy computation. By minimizing Eq. 2, we can maximize the conditional entropy based on the classifier $f(.)$. When the output of the classifier for all categories is uniform across all categories, the conditional entropy reaches its maximum, thereby reducing the semantic information contained in $f(z_{d})$.
Methods daytime-clear night-sunny dusk-rainy night-rainy daytime-foggy Average
Faster R-CNN [33] 54.7 34.0 30.5 14.0 32.2 27.7
SW [32] 50.6 33.4 26.3 13.7 30.8 26.1
IBN-Net [31] 49.7 32.1 26.1 14.3 29.6 25.5
IterNorm [14] 43.9 29.6 22.8 12.6 28.4 23.4
ISW [7] 51.3 33.2 25.9 14.1 31.8 26.3
S-DGOD [41] 56.1 36.6 28.2 16.6 33.5 28.8
CLIP-Gap [60] 51.3 36.9 32.3 18.7 38.5 31.6
Prompt-Driven [28] 53.6 38.5 33.7 19.2 39.1 32.6
OA-DG [58] 55.8 38.0 33.9 16.8 38.3 31.8
Ours 56.1 42.0 35.4 19.2 39.3 34.0
Table 1: Single-Domain Generalization Object Detection Results. We trained the model using daytime-clear as the source and subsequently tested it within that domain, while also assessing its generalization across the other four domains. The results presented in the table indicate that our Diversity Invariant Detection Model (DIDM) significantly enhances the model’s generalization performance, with bold text highlighting optimal performance. This underscores the adaptability and effectiveness of DIDM across various environments.
Lastly, in addition to estimating semantic information, we introduce a feature diversity (FD) loss to promote diverse characteristics among features. Specifically, let the domain-specific feature for each proposal be $z_{d}^{i}$. We calculate the cosine similarity between each pair $z_{d}^{i}$ and $z_{d}^{j}$ for all proposals. Given that each proposal may capture different information due to variations in location or domain, our goal is to increase the distinctions between $z_{d}^{i}$ and $z_{d}^{j}$, where $i,j\in[1,K]$ and $K$ is the number of proposals. Consequently, the feature diversity loss can be expressed as:
$\mathcal{L}_{\mathcal{FD}}=-\sum\limits_{i=1}^{K}log(\dfrac{e^{sim(f^{\prime}(z_{d}^{i}),f^{\prime}(z_{d}^{i}))}}{\sum\limits_{j=1}^{K}e^{sim(f^{\prime}(z_{d}^{i}),f^{\prime}(z_{d}^{j}))}})$ (3)
Here $sim(.)$ denotes the cosine similarity between feature mappings. $\hat{y}_{i}$ is set as 1 when $i=j$ and to 0 otherwise. $f^{\prime}(.)$ is the function of the fully connected layer. The diversity loss is designed to encourage a wide range of predicted categories, maximizing the dissimilarity among features.
@@
where $\lambda_{2}$ is the hyperparameter for balancing the loss, with the supervision of the DLM, our model can both limit the semantic information of the domain-specific feature, while keeping the feature diversity.
3.2.2 Weighted Aligning Module (WAM)
DLM focuses on learning domain-specific features, while WAM is introduced to promote domain invariance. Considering that overemphasis on domain invariance may influence domain diversity, we implement the loss weight within the feature alignment in WAM. Specifically, during the training phase, we align the original feature $z_{s}$ and the argument features $z_{a}$. Simultaneously, we minimizing the difference between $z_{s}$ and $z_{a}$, while maximizing the distance between $z_{s}$, $z_{a}$ and $z_{d}$ for the alignment:
$\mathcal{L}_{fa}$ $=1-\mathcal{L}(z_{s},z_{a})$ (5)
$\mathcal{L}_{fs}=$ $\dfrac{1}{2}(\mathcal{L}(z_{s},z_{d})+\mathcal{L}(z_{a},z_{d}))$ (6)
$\mathcal{L}(z_{s},z_{a})=$ $\dfrac{sim(z_{s}^{i},z_{a}^{i})}{sim(z_{s}^{i},z_{a}^{i})+\sum\limits_{j\neq i,j=1}^{K}sim(z_{s}^{i},z_{a}^{j})}$ (7)
By aligning features across different domains, we encourage the model to learn domain-invariant features. However, solely constraining domain disparity can reduce feature diversity. To prevent the model from overly restricting domain-specific information, we introduce a weighted parameter for feature alignment. When detection results based on $z_{s}$ and $z_{a}$ are similar, the differences between $z_{s}$ and $z_{a}$ have minimal impact, suggesting that the alignment loss should be down-weighted to preserve feature diversity. Specifically, the weight for the alignment loss is defined as:
@@ 3.2.2 Weighted Aligning Module (WAM)
where $\mathcal{L}_{c}=\sum\limits_{i=1}^{K}KL(p_{s}^{i},p_{a}^{i})$, and $\mathcal{L}_{b}=||b_{s}^{i}-b_{a}^{i}||_{2}^{2}$ from [12]. $p_{s}^{i},p_{a}^{i}$ are the classification results of $z_{s}$ and $z_{a}$, respectively, while $b_{s}^{i}-b_{a}^{i}$ are the bounding box regression results. Finally, the loss function of the weighted alignment module (WAM) is defined as follows:
$\mathcal{L}_{\mathcal{WAM}}=\beta\mathcal{L}_{fa}+\mathcal{L}_{fs}$ (9)
By training with the loss weight $\beta$, the model down weight the alignment loss when the differences between the original and argument features do not influence the detection results. This operation prevents the overemphasis on the feature alignment. Thus, the WAM can align the feature without compromising the feature diversity. Thus, the feature learning of the model is further enhanced.
3.2.3 Overall Optimization Objective
The overall optimization objective of the model is the following:
$\mathcal{L}_{total}=\mathcal{L}_{det}+\alpha\mathcal{L}_{\mathcal{DLM}}+\gamma\mathcal{L}_{\mathcal{WAM}}$ (10)
@@ 3.2.3 Overall Optimization Objective
Here $\mathcal{L}_{det}=\mathcal{L}_{reg}+\mathcal{L}_{cls}$ is the detection loss, $\alpha$ and $\gamma$ is the hyperparameters for balancing the loss. With the co-training of the detection loss, DLM and WMA, our model achieve effective domain generalization on the task.
Figure 3: Qualitative evaluation results for night-sunny and night-rainy weather conditions are presented. The top row showcases the detection visualizations generated by the Faster R-CNN [33] model, while the bottom row displays the detection visualizations produced by the DIDM. The images on the left depict night-sunny conditions, whereas those on the right illustrate night-rainy conditions.
4 Experiments
4.1 Experimental Setup
@@
Datasets. To verify the validity of the method, we utilize the datasets created by [41]. Five distinct weather conditions are included: daytime-clear, night-sunny, dusk-rainy, night-rainy, and daytime-foggy. These datasets are primarily derived from 27,708 images captured during daytime-clear conditions and 26,158 images taken at night-sunny conditions, collected from BDD-100K [45]. Additionally, 3,775 images with foggy conditions were sourced from Adverse Weather [15] and Cityscape datasets [5]. The datasets with rain rendered using [43], including 3,501 images from dusk-rainy days and 2,494 images from rainy nights. In this paper, we employ 19,395 daytime images of sunny conditions as the training set to train the model, while another 8,313 images are used to evaluate the model’s performance. The remaining four datasets are treated as unseen target domains to assess the model’s generalization capabilities. In these datasets, we focus on seven common categories: car, bike, bus, rider, person, motor, and truck.
@@
Ours 43.5 40.1 65.1 22.4 45.2 32.5 45.3 42.0
Table 2: The quantitative results (%) on the night-sunny.
Implementation Details. We utilize Faster R-CNN [33] as a two-stage detector, with ResNet101 [11] serving as the backbone network. The base network employs weights that have been pre-trained on the ImageNet dataset [9]. To train the model, we apply the stochastic gradient descent (SGD) algorithm with a momentum parameter of 0.9. The initial learning rate is set to 0.02 and decays every three epochs, while the batch-size is 4. All reported mean average precision (mAP) values are based on an Intersection over Union (IoU) threshold of 0.5.
Data Augmentation Setting. To enhance the diversity of the source domain, we employ techniques such as random clipping, the addition of Gaussian noise, grayscale enhancement, and color transformation to expand the dataset.
4.2 Performance Analysis
The DIDM we studied is compared here with four feature-based normalization methods, they are SW [32], IBN-Net [31], IterNorm [14] and ISW [7]. Comparisons were also made with the latest S-DGOD [41], CLIP-Gap [60], Prompt-Driven [28], and OA-DG [58]methods.
Figure 4: Qualitative evaluation results for night-sunny and night-rainy weather conditions are presented. The top row showcases the detection visualizations generated by the Faster R-CNN [33] model, while the bottom row displays the detection visualizations produced by the DIDM. The images on the left depict daytime-foggy conditions, whereas those on the right illustrate dusk-rainy conditions.
@@ 4.2 Performance Analysis
Ours 31.6 12.1 38.3 3.8 12.8 10.6 25.0 19.2
Table 3: The quantitative results (%) on the night-rainy.
Results on all datasets. Tab. 1 presents the testing and generalization results of DIDM on the DWD datasets [41]. In this study, we utilized the daytime-clear dataset to train the model, tested it on the same dataset, and evaluated its generalization capabilities on four additional datasets: night-sunny, dusk-rainy, night-rainy, and daytime-foggy. As shown in Tab. 1, our approach achieves the highest average generalization performance of 34.0%. When compared to the baseline network, Faster R-CNN [33], our method demonstrates improvements of 7.1% and 4.9% on the daytime-foggy and dusk-rainy datasets, respectively. Additionally, performance on the night-sunny dataset was significantly enhanced by 8.0%, while the night-rainy dataset, characterized by a complex environment, saw an improvement of 5.2%.
@@
FR [33] 22.9 27.6 55.8 29.6 33.1 34.9 21.3 32.2
@@
Ours 38.5 31.6 62.1 35.8 36.8 42.7 27.3 39.3
Table 4: The quantitative results (%) on the daytime-foggy.
Results on night-sunny conditions. As shown in the results presented in Tab. 2, our method exhibits outstanding performance across several object detection categories, particularly in the Bike, Car, and Person categories. It achieves significant performance improvements compared to all methods in the table. In these categories, our method attains detection accuracies of 40.1%, 65.1%, and 45.2%, respectively. This demonstrates the advantages of our approach in feature normalization and target detection. In addition, we achieve an overall performance improvement of 5.4%, 5.1%, and 3.5% compared to the latest S-DGOD [41], CLIP-Gap [60] and prompt-D [28], respectively. This further demonstrates the superior adaptability and robustness of our method in handling data across different domains. Fig. 3 further confirms that the detection accuracy of our method surpasses that of the baseline Faster R-CNN [33].
Results on night-rainy conditions. The night-rainy presents a particularly challenging situation. Low visibility and complex lighting significantly impacted the color-dependent model, while the rain further diminished visual clarity. Tab. 3 illustrates that our method significantly outperforms other approaches in automobile category detection, particularly in bus detection, with improvements of 7.2%, 3%, and 6% compared to S-DGOD [41], CLIP-Gap [60] and Prompt-D [28], respectively. Fig. 3 illustrates that our method can recognize objects more accurately in harsh environments, highlighting its advantages in enhancing the robustness of object detection.
@@
Ours 41.6 26.3 66.6 16.6 30.9 21.9 44.1 35.4
Table 5: The quantitative results (%) on the dusk-rainy.
@@
Results on daytime-foggy conditions. As shown in Tab. 4, our method demonstrates superior generalization performance under daytime-foggy conditions. Additionally, Fig. 4 illustrates that our method achieves higher accuracy in detecting small figure targets in foggy environments compared to Faster R-CNN [33].

### v2

@@ III-C Diversity Learning Module (DLM)
Specifically, the Diversity Learning Module (DLM) operates as follows: first, to capture the variation of the augmented features relative to the original features, we draw inspiration from OCR [47] and define the relationship between the original and augmented features as $z_{a}=\lambda_{1}z_{s}+(1-\lambda_{1})z_{d}$ where $\lambda_{1}$ is a gradually increasing hyperparameter, reflecting the model’s improving capacity to learn domain-invariant features over the process of training. The difference between $z_{s}$ and $z_{a}$ diminishes as the model becomes less sensitive to domain-specific features. Ultimately, this approach enables us to obtain the domain-specific feature $z_{d}$ during training.
@@
where $C$ represents the number of categories, $f(z_{d})\in R^{K\times C}$ is prediction function for $z_{d}$. Inspired by [47], to constrain the category semantic information in $z_{d}$, we implement to maximization of the entropy of the prediction results from $z_{d}$.
@@
where $\mathcal{H}$ represents entropy computation. By minimizing Eq. 7, we can maximize the conditional entropy based on the classifier $f(.)$. When the output of the classifier for all categories is uniform across all categories, the conditional entropy reaches its maximum, thereby restricts the semantic information contained in $f(z_{d})$.
Lastly, in addition to limiting semantic information, we introduce a feature diversity (FD) loss to promote diverse characteristics among features. Specifically, let the domain-specific feature for each proposal be $z_{d}^{i}$. We calculate the cosine similarity between each pair $z_{d}^{i}$ and $z_{d}^{j}$ for all proposals. Given that each proposal may capture different information due to variations in location or domain, our goal is to increase the distinctions between $z_{d}^{i}$ and $z_{d}^{j}$, where $i,j\in[1,K]$ and $K$ is the number of proposals. Consequently, the feature diversity loss can be expressed as:
$\mathcal{L}_{\mathcal{FD}}=-\sum\limits_{i=1}^{K}log(\dfrac{e^{(sim(f^{\prime}(z_{d}^{i}),f^{\prime}(z_{d}^{i}))/\tau)}}{\sum\limits_{j=1}^{K}e^{(sim(f^{\prime}(z_{d}^{i}),f^{\prime}(z_{d}^{j}))/\tau)}})$ (8)
Where $\tau$ denotes the temperature hyperparameter, which is utilized to control for differences between features and to encourage the model to achieve greater discrimination between various domain-specific features $z_{d}$. Here $sim(.)$ denotes the cosine similarity between feature mappings. $\hat{y}_{i}$ is set as 1 when $i=j$ and to 0 otherwise. $f^{\prime}(.)$ is the function of the normalization layer. The diversity loss is designed to encourage a wide range of predicted categories, maximizing the dissimilarity among features.
@@
Where $\lambda_{2}$ is the hyperparameter for balancing the loss. Under the supervision of the DLM, our model minimizes the semantic information in domain-specific features while maintaining their diversity, thereby enhancing the diversity of source domain distributions during training and improving the model’s robustness to unseen target domains.
III-D Weighted Aligning Module (WAM)
In addition to preserving feature diversity through the Diversity Learning Module (DLM), we propose the Weighted Alignment Module (WAM) to learn domain-invariant features while maintaining feature diversity. Traditional feature alignment methods focus solely on addressing domain discrepancy during domain alignment, which may lead to an over-tight domain-invariant feature space. This kind of distribution over-smoothing compromises feature diversity—a critical factor for generalization performance on the unseen target domain. Therefore, our goal is not to maximize alignment but to achieve a dynamic trade-off between diversity and alignment. To this end, the proposed WAM dynamically adjusts the alignment loss strength via a variable weight parameter.
Specifically, during the training phase, we align the original feature $z_{s}$ and the argument features $z_{a}$:
$\mathcal{L}(z_{s},z_{a})=$ $-\dfrac{sim(z_{s}^{i},z_{a}^{i})}{sim(z_{s}^{i},z_{a}^{i})+\sum\limits_{j\neq i,j=1}^{K}sim(z_{s}^{i},z_{a}^{j})}$ (10)
By aligning features across different domains, we encourage the model to learn domain-invariant features. However, solely constraining domain disparity can reduce feature diversity. To achieve effective feature alignment while preserving feature diversity, we introduce a weight parameter that dynamically adjusts the weight of the feature alignment loss based on the similarity between the detection results of $z_{s}$ and $z_{a}$. Specifically, when detection results based on $z_{s}$ and $z_{a}$ are similar, the differences between $z_{s}$ and $z_{a}$ have minimal impact, suggesting that the alignment loss should be down-weighted to preserve feature diversity. The weight for the alignment loss is defined as:
@@ III-D Weighted Aligning Module (WAM)
[…]
$\mathcal{L}_{\mathcal{WAM}}=\beta(1+\mathcal{L}(z_{s},z_{a}))$ (12)
By training with the weight parameters $\beta$, the model decreases the weight of the alignment loss when the distinctions between the original and augmented features do not influence the detection results. This dynamic adjustment mechanism enables the model to balance the relationship between feature alignment and feature diversity throughout the training process. Consequently, the WAM further enhances the model’s ability to learn features.
III-E Overall Optimization Objective
Overall optimization objective of the model is as follows:
$\mathcal{L}_{total}=\mathcal{L}_{det}+\alpha(\mathcal{L}_{\mathcal{DLM}}+\mathcal{L}_{\mathcal{WAM}})$ (13)
@@ III-E Overall Optimization Objective
Here $\mathcal{L}_{det}=\mathcal{L}_{reg}+\mathcal{L}_{cls}$ is the detection loss, $\alpha$ is the hyperparameters for balancing the loss. With the co-training of the detection loss, DLM and WMA, our model achieve effective domain generalization on the task.
TABLE I: Single-Domain Generalization Object Detection Results. The Results Presented in the Table Indicate that Our Diversity Invariant Detection Model (DIDM) Significantly Enhances the Model’s Generalization Performance, with Bold Text Highlighting Optimal Performance. This Underscores the Effectiveness of DIDM Across Various Environments.
Methods Ref Daytime - Clear Night - Sunny Dusk - Rainy Night - Rainy Daytime - Foggy Average
Faster R-CNN [1] NeurIPS’15 57.3 38.1 32.3 14.6 34.5 29.9
SW [48] ICCV’19 50.6 33.4 26.3 13.7 30.8 26.1
IBN-Net [42] ECCV’18 49.7 32.1 26.1 14.3 29.6 25.5
IterNorm [44] CVPR’19 43.9 29.6 22.8 12.6 28.4 23.4
ISW [43] CVPR’21 51.3 33.2 25.9 14.1 31.8 26.3
SHADE [49] CVPR’22 - 36.6 29.5 16.8 33.9 28.4
S-DGOD [45] CVPR’22 56.1 36.6 28.2 16.6 33.5 28.8
CLIP-Gap [39] CVPR’23 51.3 36.9 32.3 18.7 38.5 31.6
SRCD [50] TNNLS’24 - 36.7 28.8 17.0 35.9 29.6
Prompt-Driven [15] CVPR’24 53.6 38.5 33.7 19.2 39.1 32.6
Diversification [40] CVPR’24 50.6 39.4 37.0 22.0 35.6 33.5
DivAlign [40] CVPR’24 52.8 42.5 38.1 24.1 37.2 35.5
OA-DG [41] AAAI’24 55.8 38.0 33.9 16.8 38.3 31.8
UFR [14] CVPR’24 58.6 40.8 33.2 19.2 39.6 33.2
Sdg-yolov8 [51] Elsevier’25 - 41.1 39.0 24.6 36.0 35.2
DRSF [52] arxiv’25 57.7 40.2 39.8 20.2 38.6 34.7
TGFD[53] Elsevier’25 52.4 36.8 31.3 16.7 38.2 30.8
Ours - 57.7 43.4 36.3 20.4 40.1 35.1
IV Experiments
In this section, we conduct several experiments on the proposed DIDM model with different transfer tasks, including DWD datasets, Real to Artistic task, Diversified Weather task, and Artificial to Real task.
IV-A Experimental Setup
Implementation Details. We utilize Faster R-CNN [1] as a base detector, with ResNet101 [30] backbone. The ImageNet [54] pretrained weights are employed for initialization. To train the model, we apply the stochastic gradient descent (SGD) algorithm with a momentum parameter of 0.9. The initial learning rate is set to 0.02 and decays every three epochs. All reported mean average precision (mAP) values are based on an Intersection over Union (IoU) threshold of 0.5.
Data Augmentation Setting. To further enhance the generalization of the model to the source domain data, a series of data augmentation techniques are employed to expand the source domain dataset. These techniques include random clipping, greyscale enhancement, and color transformation, which aim to increase the model’s resilience to color variations. These augmentation methods significantly enrich the diversity of the dataset, allowing the model to be exposed to a wider range of features during the training process, thereby demonstrating improved adaptability and generalization capabilities in practical applications.
IV-B Datasets
@@
DWD datasets are included: Daytime - Clear, Night - Sunny, Dusk - Rainy, Night - Rainy, Daytime - Foggy. These datasets are primarily derived from 27,708 images captured during daytime-clear conditions and 26,158 images taken at night-sunny conditions, collected from BDD-100K [55]. Additionally, 3,775 images with foggy conditions were sourced from Adverse Weather [56] and Cityscape datasets [57]. The datasets with rain rendered using [58], including 3,501 images from dusk-rainy days and 2,494 images from rainy nights. In this paper, we employ 19,395 daytime images of sunny conditions as the training set to train the model, while another 8,313 images are used to evaluate the model’s performance. The remaining four datasets are treated as unseen target domains to assess the model’s generalization capabilities. In these datasets, we focus on seven common categories: car, bike, bus, rider, person, motor, and truck.
Real to Artistic comprises four distinct types of datasets: Pascal VOC [59], Clipart1k [60], Watercolor2k [60], and Comic2k [60]. In this study, we utilize the Pascal VOC 2007 and VOC 2012 datasets, which consist of real images, to train the model. Subsequently, we then test the model on three additional datasets to evaluate its generalization performance on art-style images. Specifically, the Clipart1k dataset includes 20 categories that overlap with those in Pascal VOC, featuring a total of 1,000 clipart-style images. In contrast, Watercolor2k and Comic2k each contain 6 categories, representing subsets of the Pascal VOC categories, with Watercolor2k focusing on watercolor styles and Comic2k on manga styles.
Diversified Weather comprises datasets from four distinct environments: Cityscapes [57], Foggy Cityscapes[61], Rainy Cityscapes [62], and BDD-100K [55]. The Cityscapes datasets offer high-quality imagery of urban street scenes, while Foggy Cityscapes, Rainy Cityscapes, and BDD-100K further enhance the model’s capability to comprehend various environments. In this study, we train the model using one of the datasets and subsequently evaluate its performance with the remaining datasets to assess its adaptability and robustness in unfamiliar environments.
Artificial to Real is a collection of synthetic images and two datasets from different environments: Sim10k [63], Cityscapes [57], and BDD-100K [55]. SIM10k [63] consists of 10,000 artificial images rendered by CTA-V, which serve as the source domain for training the model. The model is then tested on the other datasets, demonstrating detection results in the car category to validate the feasibility of its cross-domain performance.
TABLE II: The Quantitative Results (%) on the Night-Sunny.
@@
UFR [14] 43.6 38.1 66.1 14.7 49.1 26.4 47.5 40.8
Ours 43.8 40.0 66.1 25.2 46.5 36.1 46.3 43.4
IV-C Performance on DWD datasets
The DIDM we studied is compared with four feature-based normalization methods they are SW [48], IBN-Net [42], IterNorm [44] and ISW [43]. Comparisons were also made with the latest SHADE [49], S-DGOD [45], CLIP-Gap [39], SRCD [50], Prompt-Driven [15], DivAlign [40], OA-DG [41], UFR [14], SDG [51], DRSF [52] and TGFD[53] methods.
Results on DWD datasets. Table I presents the testing and generalization results of the DIDM on the DWD datasets [45]. In this study, we utilized the Daytime-Clear datasets to train the model, tested it on the same datasets, and evaluated its generalization capabilities on four additional datasets: Night-Sunny, Dusk-Rainy, Night-Rainy, and Daytime-Foggy. As shown in Table I, our approach achieves an average generalization performance of 35.0% and performs well in several scenarios, particularly under Night - Sunny and Daytime - Foggy conditions, where it attains the best performance metrics of 43.2% and 40.1%, respectively. When compared to the baseline network, Faster R-CNN [1], our method demonstrates improvements of 5.6% and 4.0% on the Daytime-Foggy and Dusk-Rainy datasets, respectively. Additionally, performance on the Night-Sunny datasets was significantly enhanced by 5.1%, while the Night-Rainy datasets, characterized by a complex environment, saw an improvement of 5.8%. This demonstrates the robustness and adaptability of our method across different environments, particularly in challenging scenarios. The performance increase is caused by enhancing the feature distribution enhancement with feature diversity, where the distance between the augmented source domain and the target domain is reduced.
TABLE III: The Quantitative Results (%) on the Night - Rainy.
@@ IV-C Performance on DWD datasets
UFR [14] 29.9 11.8 36.1 9.4 13.1 10.5 23.3 19.2
Ours 33.8 19.2 41.2 1.2 11.6 11.2 24.7 20.4
TABLE IV: The Quantitative Results (%) on the Daytime - Clear.
@@
FR [1] 59.5 47.6 76.7 49.3 56.0 51.4 60.3 57.3
SW [48] 62.3 42.9 53.3 49.9 39.2 46.2 60.6 50.6
IBN-Net [42] 63.6 40.7 53.2 45.9 38.6 45.3 60.7 49.7
IterNorm [44] 58.4 34.2 42.4 44.1 31.6 40.8 55.5 43.9
ISW [43] 62.9 44.6 53.5 49.2 39.9 48.3 60.9 51.3
S-DGOD [45] 68.8 50.9 53.9 56.2 41.8 52.4 68.7 56.1
CLIP-Gap [39] 55.0 47.8 67.5 46.7 49.4 46.7 54.7 52.5
Prompt-D [15] - - - - - - 53.6
UFR [14] 66.8 51.0 70.6 55.8 49.8 48.5 67.4 58.6
Ours 60.5 48.6 76.3 49.3 56.0 53.2 60.4 57.7
Results on Night-Sunny conditions. As shown in the results presented in Table II, our method exhibits outstanding performance across several object detection categories, particularly in the Bus, Car, and bike categories, our method attains detection accuracies of 43.8%, 66.1%, and 40.0%, respectively. This demonstrates the advantages of our approach in feature normalization and target detection. In addition, we achieve an overall performance improvement of 6.5%, 6.7%, 4.9%, and 2.6% compared to the latest CLIP-Gap [39], SRCD [50], prompt-D [15] and UFR [14], respectively. This further demonstrates the superior adaptability and robustness of our method in handling data across different domains.
Figure 3: Qualitative evaluation results for night-sunny, night-rainy daytime-foggy and dusk-rainy weather conditions are presented. The images, arranged from top to bottom, display the visualization results generated by Faster R-CNN [1], CLIP-Gap [39], DIDM, and the Ground Truth, respectively.
TABLE V: The Quantitative Results (%) on the Daytime - Foggy.
Methods Bus Bike Car Motor Person Rider Truck mAP
FR [1] 32.0 29.1 56.7 28.7 34.1 38.8 22.1 34.5
@@
UFR [14] 36.9 35.8 61.7 33.7 39.5 42.2 27.5 39.6
Ours 41.1 33.2 62.5 33.5 37.2 44.7 28.7 40.1
Results on Night-Rainy conditions. The Night - Rainy presents a particularly challenging situation. Low visibility and complex lighting significantly impacted the color-dependent model, while the rain further diminished visual clarity. Table III illustrates that our method significantly outperforms other approaches in automobile category detection, particularly in bus detection, with improvements of 5.2%, 7.3%, 8.2% and 3.9% compared to CLIP-Gap [39], SRCD [50], prompt-D [15] and UFR [14], respectively.
TABLE VI: The Quantitative Results (%) on the Dusk - Rainy.
@@
UFR [14] 37.1 21.8 67.9 16.4 27.4 17.9 43.9 33.2
Ours 41.9 30.1 68.2 18.8 29.9 19.9 45.2 36.3
Results on Daytime-Clear conditions. We evaluate the proposed method in this paper using a dataset that belongs to the same domain as the training set, as illustrated in Table IV. The results indicate that the proposed method performs exceptionally well across several detection categories, particularly in the Car, Person, and Rider categories. It shows a distinct advantage over other methods, achieving of 76.3%, 56.0%, and 53.2%, respectively. Compared to the baseline Faster R-CNN [1] method, this approach demonstrates notable improvements, thereby validating its effectiveness in the target detection task.
@@
Results on Daytime-Foggy conditions. As shown in Table V, our method demonstrates superior generalization performance under daytime-foggy conditions, achieving 40.1% of the best detection performance compared to all methods listed in the table.